michael is 17 years older than John. In 4 years the sun of their ages will be 49. How old are they now
Answer by macston(5194)

M=Michael's age: J=Johns age=M-17 yrs
If in 4 years the sum will be 49 years, currently the sum is 41 years so:
M+J=41 years Substitute for J.
M+M-17yrs=41yrs Add 17 yrs to each side.
2M=58 yrs Divide each side by 2.
M=29 yrs ANSWER 1: Michael is 29 years old.
J=M-17yrs=29yrs-17yrs=12 yrs ANSWER 2: John is 12 years old.
CHECK:
49 yrs=(M+4yrs)+(J+4yrs)
49 yrs=29yrs+4yrs+12yrs+4yrs
49 yrs=49 yrs